# Quantum memory assisted observable estimation

### v3

@@ 4 Summary
In summary, we propose a new method (CPS) to estimate the expectation values of multi-qubit observables. The method uses the QSP technique to encode information from a multi-qubit processor into a single qubit quantum memory, which allows to overcome the shot noise bottleneck of the conventional QEE. Compared to the QEE, the CPS obtains an Heisenberg limited scaling of the estimation variance with the number of Pauli strings in the decomposition of the observable. This scaling represents an improvement of $\leavevmode\nobreak\ 1/N$ compared to the QEE. We note that this improvement has been estimated assuming that there is on the order of $N$ non-commuting Pauli strings in the observable decomposition. If there are commuting sets of Pauli strings they can in principle be measured in parallel using the QEE while the CPS does not straightforwardly support parallel encoding of commuting Pauli strings. Thus, we imagine that potential trade-offs between commuting sets and non-commuting set of Pauli strings can be made, resulting in optimal strategies consisting of both methods depending on the specific observable.
@@
L.M. was supported by the Netherlands Organisation for Scientific Research (NWO/OCW), as part of the Quantum Software Consortium program (project number 024.003.037 / 3368). J.B. acknowledge funding from the NWO Gravitation Program Quantum Software Consortium.
@@ Appendix A CPS method
Hence, the success probability of the method is at least $1-2\varepsilon_{QSP}$. We select $h(\theta_{j})=\tau_{j}\cos{(\pm\theta_{j})}$, $\tau_{j}=s_{j}\epsilon|a_{j}|$, $\epsilon\in(0,1)$, $s_{j}=\text{sign}({a_{j}\langle P_{j}\rangle)}$ and the input is $\theta_{j}=\arccos{(|\langle P_{j}\rangle|)}$. Then for every $\theta_{j}$, we can write
@@
where $J_{m}(\tau)$ is the $m$-th Bessel function of the first kind. The Bessel function $J_{m}(\tau)$ is bounded, for real $\tau$ and integer $n$ as
@@
Under the conditions $|\tau|\leq k$ and $l!>(l/e)^{l}$ we can introduce the upper bound
$2\sum_{m=k+1}^{\infty}|J_{m}(\tau)|\leq 4\sum_{m=k+1}^{\infty}\frac{1}{|m|!}\Big{|}\frac{\tau}{2}\Big{|}^{|m|}<\frac{4}{(k+1)!}\Big{|}\frac{\tau}{2}\Big{|}^{k+1}<4\Big{|}\frac{e\tau}{2(k+1)}\Big{|}^{k+1}.$ (14)
Thus, we can truncate the series (12) at some point $k>n/2$ and keep the first $k$ terms. The error of the approximation of every $\exp{(i\tau_{j}\cos{[\arccos{(|\langle P_{j}\rangle|)}]})}$, $j=1,\dots,N$ is scaled super-exponentially as
$\varepsilon_{QSP}\leq O\left(\left(\frac{e\tau}{2(k+1)}\right)^{k+1}\right),$ (15)
where
$|\tau|\leq k=n/2,$ (16)
holds. Taking $k=e(\tau+\gamma)/2$, where $\gamma>0$ and $\tau=O(\gamma)$ we get the following amount of quarries [21]:
$n_{QSP}\equiv n=O\left(\frac{\log{(1/\varepsilon_{QSP})}}{\log{\log{(1/\varepsilon_{QSP})}}}\right).$ (17)
If we select |$\ket{+}_{m}\otimes\ket{\Psi_{0}}$ as the initial state the action of the ideal QSP unitary $\mathcal{U}^{{\rm ideal}}_{P_{j}}$ on this state is the following
@@
On the other hand, choosing the initial state of the ancillary qubit as |$\ket{-}_{m}$, we obtain
@@
In this case the functional transformation we select is $A(\theta)-iC(\theta)$. Hence, setting the initial state of the memory ancilla as
@@
we have the following action
@@
If we could approximate $h(\theta)$ perfectly, the phase encoded in the memory qubit would be the following
$\Theta_{j}\equiv\tau_{j}|\langle P_{j}\rangle|.$ (22)
@@
However, since QSP method is just approximating the function from the random variable, we have to take into account the error emerging in the QSP approximation:
$\Big{|}f_{QSP}\left(\arccos{(|\langle P_{j}\rangle|)}\right)-\exp{\left(i\Theta_{j}\right)}\Big{|}<\varepsilon_{QSP},$ (23)
where $f_{QSP}$ is the approximation of $\exp{\left(i\tau_{j}\cos{(\cdot)}\right)}$. Then we can write
$f_{QSP}\left(\arccos{(|\langle P_{j}\rangle|)}\right)=\exp{(i\Theta_{j})}\pm r_{j}\exp{(i\delta_{j})},$ (24)
where $|r_{j}\exp{(i\delta_{j})}|<\varepsilon_{QSP}$, $\delta_{j}\in[0,2\pi)$, $r_{j}\in R$ is an error inserted by the QSP approximation.
Let us estimate the effect of this error on the variance of the estimate of $O$. We consider the input state $(\alpha\ket{+}_{m}+\beta\ket{-}_{m})\ket{\Psi_{0}}$. After $N$ full QSP rounds, we get the final state stored in the memory qubit:
$|{\Phi}_{N}\rangle$ $=$ $\Bigg{(}(e^{i\Theta_{1}}\pm r_{1}e^{i\delta_{1}})(e^{i\Theta_{2}}\pm r_{2}e^{i\delta_{2}})\dots(e^{i\Theta_{N}}\pm r_{N}e^{i\delta_{N}})\alpha\ket{+}_{m}$
$+$ $(e^{-i\Theta_{1}}\pm r_{1}e^{-i\delta_{1}})(e^{-i\Theta_{2}}\pm r_{2}e^{-i\delta_{2}})\dots(e^{-i\Theta_{N}}\pm r_{N}e^{-i\delta_{N}})\beta\ket{-}_{m}\Bigg{)}|\theta_{1}^{N}\rangle,$
where we used the notation
$|\theta_{1}^{N}\rangle\equiv(|\theta_{1}^{+}\rangle+e^{-in\theta_{1}}|\theta_{1}^{-}\rangle)(|\theta_{2}^{+}\rangle+e^{-in\theta_{2}}|\theta_{2}^{-}\rangle)\dots(|\theta_{N}^{+}\rangle+e^{-in\theta_{N}}|\theta_{N}^{-}\rangle),$ (26)
where $n$ scales as (17). Let us for simplicity select all signs to be plus, then we can rewrite the latter state as follows
$|{\Phi}_{N}\rangle$ $=$ $\Bigg{(}\left(e^{i{\Phi}_{N}}+A\right)\alpha\ket{+}_{m}+\left(e^{-i{\Phi}_{N}}+B\right)\beta\ket{-}_{m}\Bigg{)}|\theta_{1}^{N}\rangle,$ (27)
where we used the notations ${\Phi}_{N}=\sum\limits_{j=1}^{N}\Theta_{j}$ and
$A\equiv\sum\limits_{j=1}^{N}r_{j}e^{i\delta_{j}}e^{i\sum\limits_{k=1,k\neq j}^{N}\Theta_{k}}+\sum_{\begin{subarray}{c}j,k=1,\\ j\neq k\end{subarray}}^{N}r_{j}r_{k}e^{i(\delta_{j}+\delta_{k})}e^{i\sum\limits_{r=1,r\neq j,k}^{N}\Theta_{r}}+\dots+(r_{1}r_{2}\dots r_{n})e^{i\sum\limits_{j=1}^{N}\delta_{j}},$ (30)
and $B$ is the same expression but with the minus sign in the exponential functions. We rewrite the state in Eq. (27) as
$|{\Phi}_{N}\rangle=e^{\Phi_{N}}\left(\alpha\ket{+}_{m}+\beta e^{-2i\Phi_{N}}\ket{-}_{m}\right)\otimes|\theta_{1}^{N}\rangle+(A\alpha\ket{+}_{m}+B\beta\ket{-}_{m})\otimes|\theta_{1}^{N}\rangle,$ (31)
and in the main text of the article we denote it as
@@
$\left({\alpha}|+\rangle_{m}+{\beta}e^{-2i\sum_{j=1}^{N}\tau_{j}\cos{\theta_{j}}}|-\rangle_{m}\right)\otimes|\theta_{1}^{N}\rangle+\varepsilon_{QSP}(n)\ket{\xi},$ (32)
where the second term is arising from the imperfection of the QSP approximation process. The notations we use are $\theta_{j}\equiv\arccos{(|\langle\hat{P}_{j}\rangle|)}$, $\tau_{j}\equiv\text{sign}{(a_{j}\langle\hat{P}_{j}\rangle)}\epsilon|a_{j}|$.
For the further calculation of the variance, where we are interested in the rate in $N$, we assume, for simplicity, $r_{j}=r$ and $\delta_{j}=\delta$, hold for all $j=1,\dots,N$. The series $A$ can be always written as
$A$ $=$ $C_{N}^{1}re^{i\delta}e^{i\sum\limits_{k=1,k\neq 1}^{N}\Theta_{k}}+C_{N}^{2}r^{2}e^{2i\delta}e^{i\sum\limits_{r=1,r\neq 1,2}^{N}\Theta_{r}}+\dots+C_{N}^{N}r^{N}e^{iN\delta}=Re^{i\Theta_{R}},$ (33)
$R$ $=$ $\sqrt{\sum\limits_{j=1}^{N}(C_{j}^{N}r^{j})^{2}+2\sum\limits_{j,k=1,j\neq k}^{N}C_{N}^{j}C_{N}^{k}r^{j}r^{k}\cos{\left(i\left((j-k)\delta+\sum\limits_{r=1,r\neq j}^{N}\Theta_{r}-\sum\limits_{r=1,r\neq k}^{N}\Theta_{r}\right)\right)}},$ (34)
$\tan{(\Theta_{R})}$ $=$ $\frac{C_{N}^{1}r\sin{(\delta+\sum\limits_{k=1,k\neq 1}^{N}\Theta_{k})}+C_{N}^{2}r^{2}\sin{(2\delta+\sum\limits_{r=1,r\neq 1,2}^{N}\Theta_{r})}+\dots+C_{N}^{N}r^{N}\sin{(N\delta)}}{C_{N}^{1}r\cos{(\delta+\sum\limits_{k=1,k\neq 1}^{N}\Theta_{k})}+C_{N}^{2}r^{2}\cos{(2\delta+\sum\limits_{r=1,r\neq 1,2}^{N}\Theta_{r})}+\dots+C_{N}^{N}r^{N}\cos{(N\delta)}},$ (35)
where $C_{N}^{k}$ are the binomial coefficients, $N\geq k\geq 0$. Hence, the later expression can be bounded by
$\Big{|}Re^{i\Theta_{R}}\Big{|}<\sqrt{\sum\limits_{j=1}^{N}(C_{j}^{N}\varepsilon_{QSP}^{j})^{2}+2\sum\limits_{j,k=1,j\neq k}^{N}C_{k}^{N}C_{j}^{N}\varepsilon_{QSP}^{j+k}}=\sum\limits_{j=1}^{N}C_{j}^{N}\varepsilon_{QSP}^{j}=(1+\varepsilon_{QSP})^{N}-1\sim N\varepsilon_{QSP},$ (36)
when $N\varepsilon_{QSP}<<1$. Similar derivations can be done for $B$. After renormalization we can write the state $|{\Phi}_{N}\rangle$ as follows
$|{\Phi}_{N}\rangle$ $=$ $\frac{\left(e^{-i{\Phi}_{N}}+Re^{-i\Theta_{R}}\right)|-\rangle_{m}+\left(e^{i{\Phi}_{N}}+Re^{i\Theta_{R}}\right)|+\rangle_{m}}{\sqrt{2}\sqrt{1+R^{2}+2R\cos{(\Theta_{R}-{\Phi}_{N})}}}|\theta_{1}^{N}\rangle.$ (37)
Let introduce the notations
$ae^{i\Phi_{a}}\equiv e^{i{\Phi}_{N}}+Re^{i\Theta_{R}},$ (38)
$a^{2}=R^{2}+1+2R\cos{(\Theta_{R}-\Phi_{N})}$
$\tan{\Phi_{a}}=\frac{\sin{\Phi_{N}}+R\sin{\Theta_{R}}}{\cos{\Phi_{N}}+R\cos{\Theta_{R}}}$
and rewrite the latter state as follows
|$\ket{{\Phi}_{N}}=(\cos{{\Phi}_{a}}|0\rangle_{m}+i\sin{{\Phi}_{a}}|1\rangle_{m})|\theta_{1}^{N}\rangle.$ (39)
Hence, wee need to estimate $\Phi_{a}$ that will define the $\Phi_{N}$ estimate later. The probabilities to measure $|0\rangle$ and $|1\rangle$ are the following
$P(X=0|{\Phi}_{N})=\frac{1}{2}(1+\cos{(2{\Phi}_{a})}),\quad P(X=1|{\Phi}_{N})=\frac{1}{2}(1-\cos{(2{\Phi}_{a})}),$ (40)
@@
$P(Y=0|{\Phi}_{N})=\frac{1}{2}(1-\sin{(2{\Phi}_{a})}),\quad P(Y=1|{\Phi}_{N})=\frac{1}{2}(1+\sin{(2{\Phi}_{a})}).$ (41)
Since we don’t know the preparation circuit of the state $|{\Phi}_{N}\rangle$, we can’t use QPE to estimate the phase directly. To estimate $P(X=0|\Phi_{N})$, $P(Y=0|\Phi_{N})$ and then to estimate the total accumulated phase we need to repeat all the procedure of encoding |$\ket{\Phi_{N}}$ in the QM $M_{q}$ times, every time doing the projective measurement on |$\ket{\Phi_{N}}$. Using the probability estimates, we can obtain the estimates of sine and cosine functions of ${\Phi}_{a}$.
@@
$\hat{Q}\equiv\frac{1-2\hat{P}(Y=0|{\Phi_{N}})}{2\hat{P}(X=0|{\Phi_{N}})-1}=\hat{\tan}{(2{\Phi_{a}})},$ (42)
and the variance of the estimate of ${\Phi_{a}}$ is the following
$\sigma^{2}\hat{{\Phi_{a}}}\approx\left(\frac{\partial{\Phi_{a}}}{\partial Q}\right)^{2}\sigma^{2}\hat{Q}=\left(\frac{1}{2(1+Q^{2})}\right)^{2}\sigma^{2}\hat{Q}.$ (43)

### v4

@@ 4 Summary
In summary, we propose a new method (CPS) to estimate the expectation values of multi-qubit observables. The method uses the QSP technique to encode information from a multi-qubit processor into a single qubit quantum memory, which allows to overcome the shot noise bottleneck classically summing individually estimated Pauli strings. Compared to the QEE, the CPS obtains an Heisenberg limited scaling of the estimation variance with the number of state preparations. This scaling represents an improvement of $\leavevmode\nobreak\ 1/T$. In addition, CPS achieves an improvement scaling as $\sqrt{N}$ compared to estimating each Pauli string with standard QPE followed by classical summation. We note that this improvement has been estimated assuming that there is on the order of $N$ non-commuting Pauli strings in the observable decomposition. If there are commuting sets of Pauli strings they can in principle be measured in parallel using the QEE method while the CPS method does not straightforwardly support parallel encoding of commuting Pauli strings. Thus, we imagine that potential trade-offs between commuting sets and non-commuting set of Pauli strings can be made, resulting in optimal strategies consisting of both methods depending on the specific observable. CPS can be considered as a new variation of QPE for complex observables where direct encoding of the observable as a unitary is not possible.
@@
L.M. was supported by the Netherlands Organisation for Scientific Research (NWO/OCW), as part of the Quantum Software Consortium program (project number 024.003.037 / 3368). This work has received support from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). This work is supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. J.B. acknowledge funding from the NWO Gravitation Program Quantum Software Consortium. Funded by the European Union. Views and opinions expressed are however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A CPS method
Hence, the success probability of the method is at least $1-2\varepsilon_{QSP}$.
In this paper, we are using the QSP technique to approximate $\exp{(i\sum_{j=1}^{N}\tau_{j}\cos{[\arccos{(|\langle P_{j}\rangle|)}]})}$, where we denote $\tau_{j}=s_{j}\epsilon|a_{j}|$, $s_{j}=\text{sign}({a_{j}\langle P_{j}\rangle)}$ and $a_{j}$ are the coefficients of (1). To this end, we select $h(\theta_{j})=\tau_{j}\cos{(\pm\theta_{j})}$, where the input is $\theta_{j}=\arccos{(|\langle P_{j}\rangle|)}$. Then for every $\theta_{j}$, we can write
@@
where $J_{m}(\tau)$ is the $m$-th Bessel function of the first kind. The Bessel function $J_{m}(\tau)$ is bounded, for real $\tau$ and integer $m$ as
@@
For $m\geq k$ the factorial $m!$ grows faster then the exponential $|\tau|^{m}$. As the latter decays with $m$, good approximations are obtained (13) at $m>n/2$. The sum from $m=k$ can be bounded by the first term:
$\sum_{m=k}^{\infty}\frac{1}{m!}\left(\frac{|\tau|}{2}\right)^{m}<\frac{1}{k!}\left(\frac{|\tau|}{2}\right)^{k}\cdot\sum_{m=0}^{\infty}\left(\frac{|\tau|}{2k}\right)^{m}<\frac{1}{k!}\left(\frac{|\tau|}{2}\right)^{k},$ (15)
where we used that since
$|\tau|\leq n_{QSP}/2=k,$ (16)
the ratio is $|\tau|/2k<k/2k<1$, so the geometric series converges and is less than one. Then the upper bound is [3]:
$\varepsilon_{QSP}\leq 2\sum_{m=k}^{\infty}|J_{m}(\tau)|\leq 4\sum_{m=k}^{\infty}\frac{1}{m!}\Big{|}\frac{\tau}{2}\Big{|}^{m}<\frac{4}{k!}\Big{|}\frac{\tau}{2}\Big{|}^{k}<4\left(\frac{e|\tau|}{2k}\right)^{k}.$ (17)
The last inequity is due to the Stirling’s approximation $l!>(l/e)^{l}$, for $l\geq 1$. Then we can deduce
$k=O\left(\frac{\ln{(1/\varepsilon_{QSP})}}{\ln{k}}\right).$ (18)
The logarithm in the denominator complicates direct solving. However, this suggests an iterative relation, that is a characteristic of the Lambert $W(z)$ function:
$W(z)e^{W(z)}=z.$ (19)
We attempt to rewrite our expression into the latter form. Let us introduce the notation $E=1/\varepsilon_{QSP}$. Then, we can rewrite (18) as $k\ln{k}=O(E)$. Changing the variables $y=\ln{k}$, $k=e^{y}$, we get $ye^{y}=O(E)$. This is similar to the defining equation (19) of the Lambert $W(E)=y$ function. Then $k\approx e^{W(E)}$, holds. For large values of $E$, the Lambert $W(z)$ function can be approximated as:
$W(E)=\ln E-\ln\ln E+\mathcal{O}(1).$ (20)
Then we can write
$k=e^{\ln E-\ln\ln E+\mathcal{O}(1)}=\frac{E}{\ln{E}}e^{O(1)}=O\left(\frac{E}{\ln{E}}\right)=O\left(\frac{\log{(1/\varepsilon_{QSP})}}{\log{\log{(1/\varepsilon_{QSP})}}}\right).$ (21)
The error of the approximation of every $\exp{(i\tau_{j}\cos{(\theta_{j})})}$, $j=[1,N]$ is scaled super-exponentially
$\varepsilon_{QSP}\leq O\left(\left(\frac{e\min|\tau_{j}|}{2k}\right)^{k}\right),$ (22)
and the amount of quarries, taking into account (16) and (21), is [23]:
$n_{QSP}=O\left(\max\limits_{j}{|\tau_{j}|}+\frac{\log{(1/\varepsilon_{QSP})}}{\log{\log{(1/\varepsilon_{QSP})}}}\right).$ (23)
This is the number of times the QSP circuit must repeat to approximate the target function with an error $\varepsilon_{QSP}$. We demand the standard deviation $\sqrt{\eta}$ of the estimate $\hat{O}$ to be greater than the QSP errors:
$\varepsilon_{QSP}\leq\frac{\sqrt{\eta}}{N},$ (24)
which require the number of steps in each QSP protocol to be
$n_{QSP}=O\left(\max\limits_{j}{|\tau_{j}|}+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right).$ (25)
If we select |$\ket{+}_{m}\otimes\ket{\Psi_{0}}$ as the initial state, the action of the ideal QSP unitary $\mathcal{U}^{{\rm ideal}}_{P_{j}}$ on this state is the following:
@@
On the other hand, selecting the initial state of the ancillary qubit as |$\ket{-}_{m}$, we obtain
@@
In this case, the functional transformation we select is $A(\theta)-iC(\theta)$. Hence, setting the initial state of the memory ancilla as
@@
we have the following action:
@@
Since the QSP method is just approximating the function from the random variable, we have to take into account the error accumulation in the QSP approximation. If we apply the QSP sequence on a (normalized) input state |$\ket{\psi}_{m}\otimes\ket{\Psi_{0}}$ we get the (unnormalized) state
@@
$\sim$ $\left(e^{i\tau_{j}\cos{(\theta_{j})}}{\alpha}|+\rangle_{m}+{\beta}e^{-i\tau_{j}\cos{(\theta_{j})}}|-\rangle_{m}\right)\otimes(e^{in\theta_{j}/2}\ket{\theta^{+}_{j}}+e^{-in\theta_{j}/2}\ket{\theta^{-}_{j}})+\varepsilon_{QSP}(n)\ket{\xi},$ (30)
where |$\ket{\xi}$ is a general (normalised) error state, which can be an entangled state between the memory qubit and the qubits of the multi-qubit device. We reinitialize the computational qubit by tracing it out. After $N$ rounds of such QSP encoding, with probability $(1-\varepsilon_{QSP}(n))^{N}$, $\epsilon_{QSP}(n)<<1$ the state encoded in the memory qubit is:
$|{\Phi}_{N}\rangle$ $\sim$ ${\alpha}e^{i\sum_{j=1}^{N}\tau_{j}\cos{(\theta_{j})}}|+\rangle_{m}+{\beta}e^{-i\sum_{j=1}^{N}\tau_{j}\cos{(\theta_{j})}}|-\rangle_{m}.$ (31)
With probability $(1-(1-\varepsilon_{QSP}(n))^{N})\sim N\varepsilon_{QSP}(n)$ we assume that don’t encode any reasonable information i.e. that the memory qubit is depolarized due to erroneous QSP encoding of the Pauli strings.
Selecting $\alpha=\beta$, we can rewrite the latter state as follows:
|$\ket{{\Phi}_{N}}\sim\cos{{\Phi}_{N}}|0\rangle_{m}+i\sin{{\Phi}_{N}}|1\rangle_{m},$ (32)
where the phase ${\Phi}_{N}\equiv\sum_{j=1}^{N}\tau_{j}\cos{(\theta_{j})}=\langle O\rangle$ is a target mean value. Hence, we need to estimate $\Phi_{N}$ which will provide $\langle O\rangle$. The probabilities to measure $|0\rangle$ and $|1\rangle$ are the following:
$P(X=0|{\Phi}_{N})=\frac{1}{2}(1+\cos{(2{\Phi}_{N})}),\quad P(X=1|{\Phi}_{N})=\frac{1}{2}(1-\cos{(2{\Phi}_{N})}),$ (33)
@@
$P(Y=0|{\Phi}_{N})=\frac{1}{2}(1-\sin{(2{\Phi}_{N})}),\quad P(Y=1|{\Phi}_{N})=\frac{1}{2}(1+\sin{(2{\Phi}_{N})}).$ (34)
Since we don’t know the preparation circuit of the state $|{\Phi}_{N}\rangle$, we can’t use QPE to estimate the phase directly. To estimate $P(X=0|\Phi_{N})$, $P(Y=0|\Phi_{N})$ and then to estimate the total accumulated phase, we need to repeat all the procedure of encoding |$\ket{\Phi_{N}}$ in the QM $M_{q}$ times, every time doing the projective measurement on |$\ket{\Phi_{N}}$. Using the probability estimates, we can obtain the estimates of sine and cosine functions of ${\Phi}_{N}$.
@@
${Q}\equiv\frac{1-2{P}(Y=0|{\Phi_{N}})}{2{P}(X=0|{\Phi_{N}})-1}={\tan}{(2{\Phi_{N}})}.$ (35)
The estimate of ${\Phi}_{N}$ is denoted by $\hat{\Phi}_{N}$. Then the variance of $\hat{\Phi}_{N}$ is the following:
$\sigma^{2}\hat{\Phi}_{N}\approx\left(\frac{\partial{\Phi_{N}}}{\partial Q}\right)^{2}\sigma^{2}\hat{Q}=\left(\frac{1}{2(1+Q^{2})}\right)^{2}\sigma^{2}\hat{Q}.$ (36)